# Unexpected but informative: What fixation-related potentials tell us about the processing of confusing program code

### v2

@@ Abstract
As software pervades more and more areas of our professional and personal lives, there is an ever-increasing need to maintain software, and for programmers to be able to efficiently write and understand program code. In the first study of its kind, we analyze fixation-related potentials (FRPs) to explore the online processing of program code patterns that are confusing to programmers, but not the computer (so-called atoms of confusion), and their underlying neurocognitive mechanisms in an ecologically valid setting. Relative to clean counterparts in program code without an atom of confusion, confusing code elicits a late frontal positivity of about 400 to 700 ms after first looking at the atom of confusion. This frontal positivity resembles an event-related potential (ERP) component found during natural language processing that is elicited by unexpected but plausible words in sentence context. Thus, we suggest that the brain engages similar neurocognitive mechanisms in response to unexpected and informative inputs in program code and in natural language. In both domains, these inputs update a comprehender’s situation model that is essential for information extraction from a quickly unfolding input. Our results have far-reaching implications for programming and pave the way for interdisciplinary collaborations between software engineering and psycholinguistics.
@@
Software forms an integral aspect of our everyday life. Thus, it is of high relevance to understand the processes that underlie skilled programming, the process of creation and maintenance of software. One central aspect of programming is reading and comprehending program code, which is crucial for various programming activities, such as implementing new functionalities or fixing program errors [1, 2, 3, 4, 5]. During these activities, misconceptions about the code, such as confusion about its execution, purpose, or intended message, lead to software failures with possibly fatal consequences [4, 6].
The main goal of the present study is to explore the processing of confusing elements in program code and their underlying neurocognitive mechanisms in an ecologically valid setting. Recently, various studies identified a set of small, recurring patterns [7] prevalent in program code [8, 9, 10], the so-called atoms of confusion. An atom of confusion is a syntactically correct code pattern that is unambiguous for the computer to execute, but, due to its unintuitive structure, is confusing to programmers, who encounter difficulties in understanding the code and might even misinterpret it [7, 11].
@@
R can be straightforwardly determined as 3, as several studies on atoms of confusion confirmed [12, 7, 13]. It is important to note that the confusing and clean snippet conditions are both syntactically correct and semantically equivalent, but they differ in their risk that a human programmer may misunderstand their semantics due to the atom of confusion [7]. While such subtle differences may seem insignificant at first glance, they have caused severe bugs in the past, such as a mismatching indentation in Apple’s iOS operating system rendering its secure network connections vulnerable [14, 15].
@@
Despite the insights on atoms of confusion obtained by behavioral experiments, the underlying mechanisms of the phenomenon are unclear. One missing element is our limited understanding of which neurocognitive mechanisms exactly are engaged in response to the confusing program code that is causing the difficulty. Identifying these mechanisms would allow us to more deeply understand the principles behind misconceptions in program code, with far-reaching implications for programming practice. Specifically, this line of research can provide empirical evidence for refining coding conventions, programming language design, and educational practices, ultimately leading to more secure and more reliable software.
The premise of our work is that our understanding of how humans process natural language could be the key to interpreting the processing of atoms of confusion and to discovering the underlying neurocognitive mechanisms. Indeed, parallels between programming languages and natural languages have been observed in several studies, hinting at an overlap in activated brain areas between natural language (e.g., Brodmann Areas 21, 44, and 47) and program comprehension [19, 20, 21, 22, 23]. In general, considering analogies with what is already known about human language processing and the processing of program code might prove to be a very fruitful avenue.
@@
Or should we rather think of atoms of confusion in terms of (2a) syntactic reanalysis similar to what is typically observed in garden path sentences, such as The horse raced past the barn fell, which, like an atom of confusion, in fact only has a single correct interpretation, which is however difficult to work out for a human, as there might be uncertainty about the function of the verb raced. It can either be analyzed as a main verb of the sentence or alternatively as a past participle of a reduced relative clause The horse (that was) raced past the barn fell (which is similar to indentation in our code example, where there is uncertainty about the program execution). In this case, we should expect to see a late posterior positivity (P600) [30, 31]. Besides garden path sentences that involve syntactic reanalysis, the P600 is generally observed in sentences with syntactic errors, such as The child throw the toy, but also in sentences with syntactic ambiguities, as in The woman persuaded to answer the door relative to the syntactically unambiguous sentence The woman struggled to prepare the meal [31, 32]. In these cases, the P600 is typically interpreted as a correlate of a syntactic revision (e.g., passivizing the verb persuaded and attaching it to a reduced relative clause), but it can also be evoked by semantic violations (see Paczynski and Kuperberg [33] for an example).
@@
Currently, it is unknown whether the confusion induced by atoms of confusion as perceived by programmers results (1) from semantic integration of unexpected and challenging content (in which case we would expect to see an N400), (2) from syntactic ambiguity (see below), or (3) from something else entirely (producing an unexpected ERP signal). As for syntactic ambiguity, the confusion may arise from (2a) syntactic reanalysis (in which case we would expect to see a late posterior positivity) or from (2b) integration of unexpected but plausible content into the situation model of the preceding context (in which case we would expect to see a late frontal positivity). The premise of our work is that building on natural-language studies would help us to understand the source and processing of such confusion, illuminating the intricacies of program comprehension. This, in turn, would be invaluable to better support programmers by dedicated programming languages and curricula. For example, if atoms of confusion evoke a perceived syntactic ambiguity, this could be avoided by better design choices of programming language designers (e.g., by requiring explicit brackets around bodies of for, while, and if statements); if they rather cause problems with semantic integration, this can be captured in dedicated training procedures in programming curricula and coding guidelines (e.g., defining clear and consistent rules that instruct programmers on the use of patterns). If we observe an entirely unexpected ERP signal, this could indicate that the confusion arising from the presence of an atom of confusion has no equivalent in natural-language comprehension, or that there is no unified cause for atoms of confusion. Thus, independent of the concrete results, it is highly relevant for the field of program comprehension to understand the electrophysiological signals and hence the underlying mechanisms of processing atoms of confusion. Looking at a broader picture, if we observe more and more analogies between processing program code and natural language text, this suggests that our premise of a methodological transfer between the two disciplines is worthwhile.
@@
Since it is unclear which ERP component relates to processing confusing code, we employ a data-driven approach to detect differences between confusing and corresponding clean code without targeting a hypothesized component, as research on program comprehension and its underlying neurocognitive mechanisms is still in its infancy. Contrary to research on natural language comprehension, which looks back on several decades of behavioral, neurophysiological, and brain imaging studies that have unraveled a variety of processing aspects and representational structures [37], studying program comprehension faces unique challenges, such as non-linear reading flow (e.g., program statements are processed not in lexical order but along control and data flow of possible program executions) [38, 39, 40, 41]. So far, only one study on program comprehension investigated ERP responses to syntactic and semantic errors in program code: Using a token-wise presentation of code, Kuo and Prat [42] identified P600 and N400 components for syntactically and semantically invalid program code, respectively. However, this setup of a token-wise presentation, while usually applied in reading studies with natural language materials [43], is too artificial and highly problematic for the analysis of program code: First, it enforces an unnatural linear reading flow. Second, with token-wise presentation mode, each token is presented with a uniform duration, which allows the programmer only a cursory glance at the code but does not account for calculations that can strongly vary in complexity between tokens within a code snippet. Both aspects strongly compromise the construct and ecological validity of neurocognitive studies on program comprehension.
@@
In an effort to increase the ecological validity of our study, we seek to present program code as a whole. A major challenge with presenting a snippet of code all at once, though, is that the exact point in time at which the critical segment containing the atom of confusion is processed varies both between items and between participants and, therefore, cannot be assessed with conventional ERP averaging methods. That is, simply synchronizing electroencephalographical (EEG) activity with the onset of the presentation of the code snippet leads to a temporal misalignment of the ERP activity related to the processing of the atom within the trial period and, thus, diffuses the signal. This severely reduces statistical power and interpretability. To address this key challenge, we employ a more sophisticated and powerful approach based on fixation-related potentials (FRPs). In the FRP approach, event synchronization is not realized by external stimuli [44]. Rather, eye movements and EEG are recorded simultaneously while the program code is presented as a whole, and eye-tracking data are used to assess the participants’ first eye fixations in the area of the atom of confusion (which is well-defined for every atom of confusion). This first fixation is then used for aligning the EEG signals. This way, we are able to determine the exact point in time at which information from the atoms is processed and can use it for synchronizing the electrophysiological data.
@@
To uncover which neurocognitive mechanisms exactly are engaged in response to processing confusing program code, we conducted an FRP experiment involving participants. We carefully adjusted the existing stimuli from previous work on atoms of confusion in code written in the Java programming language [13] to optimize them for an FRP setup. In particular, each snippet was presented in a condition with and without an atom of confusion (across participants). Each pair of the snippet condition had a comparable number and length of code lines. Furthermore, to select suitable fixations for FRP analysis, we ensured that each snippet contained a single, clearly defined area of interest that contains either the atom of confusion (confusing snippets) or an equivalent without atom (clean snippets) (see Section 4, for details). The experiment consisted of three blocks, as illustrated in Figure 2, in which we presented snippets (half confusing and half clean). We asked the participants to comprehend the snippets as fast and accurately as possible in order to be able to report the output. During the experiment, we collected comprehension time, answer correctness, and subjective difficulty ratings to verify that our confusing snippets were generally more difficult to process on the behavioral level than clean snippets. We collected EEG and eye-tracking data to calculate FRPs by mapping fixations to the areas of interest within the code snippets. We took special care to identify and correct eye movements and muscular artifacts by means of an independent component analysis (see Section 4, for further details). We synchronized the fixations with the EEG data stream to calculate FRPs. On the obtained FRPs, we tested for significant amplitude differences between confusing and clean code using pairwise cluster permutation tests.
@@ 2 Results
We report on the first-ever FRP analysis of program comprehension by measuring and pinpointing the processing of confusing elements and its neural underpinning during program comprehension in an ecologically valid setting. Our results show that confusing code elicits a larger frontal positivity between and after first fixating the relevant area of the atom of confusion compared to fixating clean code (without an atom). We relate the identified component to the late frontal positivity occurring in natural language processing for unexpected but plausible input which elicits an update of the situation model.
@@ 2.1 Behavioral results
The behavioral data comprised comprehension time, answer correctness, and subjective difficulty ratings, where the last two were mapped to a binary, categorical scale (the submitted answer was correct or the submitted rating was easy). The descriptive trial statistics are illustrated in the left part of Table 1.
In line with prior work, our behavioral results based on general linear mixed-effects regression (GLMER) models show that the confusing snippets took longer to comprehend, had fewer correct answers, and were rated less often as easy to understand, as displayed in the right part of Table 1. The full analysis results including all fixed and random effects are provided in the supplementary information, Appendix C.
@@ 2.2 Fixation-related potentials analysis
Our primary goal was to determine the neurocognitive processes triggered by the presence of atoms of confusion in program code and to relate those processes to ERP components typically identified during natural language comprehension. In fact, there is a large delay and large variance in the point in time when participants first fixate the critical region after onset of the snippet: the first fixation on the critical region happens, on average, after stimulus onset in the clean condition, and in the confusing condition. The stimulus-onset aligned ERPs hence cannot capture this first point in time at which information from the critical region is presumably processed. Fixation data show that on the first pass, participants fixated the critical region for in the clean condition and in the confusing condition. Therefore, we compared FRPs time-locked to the first fixation in the area of the atom of confusion and its counterpart without atom, spanning an epoch from preceding the start of the fixation until thereafter. For some trials, this epoch could not be extracted due to either bad quality in the fixation data, no fixation in the area of interest, artifacts in the EEG signal, or a too short time period until the end of the snippet viewing phase. We excluded these trials from the subsequent analysis ( of ). Then, we calculated the participants’ FRP averages for both conditions across all trials. The grand average FRP data across all participants for electrodes at frontal, central, and parietal recording sites are presented in Figure 3a.
As illustrated in Figure 3, confusing code elicits a sustained frontal positivity relative to clean code. This positivity was most pronounced between and at frontal recording sites (i.e., F3, Fz, F4). At posterior recording sites, the positive amplitude difference between confusing and clean code was attenuated in this time interval. Figure 3b shows the difference in amplitude for all electrodes in form of topographical distributions in consecutive time intervals between and . Positive differences between both conditions were also present beyond after fixation onset at some left frontal recording sites (e.g., F7 and F3), though less consistent. We statistically analyzed these effects using a cluster-based paired permutation t-test, which identifies significant clusters of adjacent data points (defined temporally and spatially).
@@
These results show that the late frontal positivity is significantly greater for confusing code than for clean code, indicating that both conditions are processed differently starting at around after onset of the first fixation.

### v3

@@ Abstract
As software pervades more and more areas of our professional and personal lives, there is an ever-increasing need to maintain software and for programmers to efficiently write and understand program code. In the first study of its kind, we analyze fixation-related potentials (FRPs) to explore the online processing of program code patterns that are confusing to programmers, but not to the computer (so-called atoms of confusion), and their underlying neurocognitive mechanisms in an ecologically valid setting. Relative to clean counterparts in program code without an atom of confusion, confusing code elicits a late frontal positivity of about 400 to 700 ms after first looking at the atom of confusion. This frontal positivity resembles an event-related potential (ERP) component found during natural language processing that is elicited by unexpected but plausible words in sentence context. Thus, we suggest that the brain engages similar neurocognitive mechanisms in response to unexpected and informative inputs in program code and in natural language. In both domains, these inputs update a comprehender’s situation model, which is essential for information extraction from a quickly unfolding input. Our results have far-reaching implications for programming and pave the way for interdisciplinary collaborations between software engineering and psycholinguistics.
@@
Software forms an integral aspect of our everyday life. Thus, it is highly relevant to understand the processes that underlie skilled programming, including the creation and maintenance of software. One central aspect of programming is reading and comprehending program code, which is crucial for various programming activities, such as implementing new functionalities or fixing program errors [1, 2, 3, 4, 5]. During these activities, misconceptions about the code, such as confusion regarding its execution, purpose, or intended message, can lead to software failures with potentially fatal consequences [4, 6].
The main goal of the present study is to explore the processing of confusing elements in program code and their underlying neurocognitive mechanisms in an ecologically valid setting. Recently, various studies have identified a set of small, recurring patterns [7] prevalent in program code [8, 9, 10], the so-called atoms of confusion. An atom of confusion is a syntactically correct code pattern that is unambiguous for the computer to execute, but, due to its unintuitive structure, is confusing to programmers, who encounter difficulties in understanding the code and might even misinterpret it [7, 11].
@@
R can be straightforwardly determined as 3, as several studies on atoms of confusion confirmed [12, 7, 13]. It is important to note that the confusing and clean snippet conditions are both syntactically correct and semantically equivalent, but they differ in their risk that a human programmer may misunderstand their semantics due to the atom of confusion [7]. While such subtle differences may seem insignificant at first glance, they have caused severe bugs in the past, such as a mismatched indentation in Apple’s iOS operating system that rendered its secure network connections vulnerable [14, 15].
@@
Despite the insights into atoms of confusion obtained from behavioral experiments, the underlying mechanisms of the phenomenon remain unclear. One missing element is our limited understanding of which neurocognitive mechanisms are engaged in response to the confusing program code that is causing the difficulty. Identifying these mechanisms would allow us to gain a deeper understanding of the principles behind misconceptions in program code, with far-reaching implications for programming practice. Specifically, this line of research can provide empirical evidence for refining coding conventions, programming language design, and educational practices, ultimately leading to more secure and reliable software.
The premise of our work is that our understanding of how humans process natural language could be the key to interpreting the processing of atoms of confusion and discovering the underlying neurocognitive mechanisms. Indeed, parallels between programming languages and natural languages have been observed in several studies, hinting at an overlap in activated brain areas between natural language (e.g., Brodmann Areas 21, 44, and 47) and program comprehension [19, 20, 21, 22, 23]. In general, considering analogies with what is already known about human language processing and the processing of program code might prove to be a very fruitful avenue.
@@
Or should we rather think of atoms of confusion in terms of (2a) syntactic reanalysis similar to what is typically observed in garden path sentences, such as The horse raced past the barn fell, which, like an atom of confusion, in fact only has a single correct interpretation that is, however, difficult to work out for a human, as there might be uncertainty about the function of the verb raced. It can either be analyzed as a main verb of the sentence or alternatively as a past participle of a reduced relative clause The horse (that was) raced past the barn fell (which is similar to indentation in our code example, where there is uncertainty about the program execution). In this case, we should expect to see a late posterior positivity (P600) [30, 31]. Besides garden path sentences that involve syntactic reanalysis, the P600 is generally observed in sentences with syntactic errors, such as The child throw the toy, but also in sentences with syntactic ambiguities, as in The woman persuaded to answer the door relative to the syntactically unambiguous sentence The woman struggled to prepare the meal [31, 32]. In these cases, the P600 is typically interpreted as a correlate of a syntactic revision (e.g., passivizing the verb persuaded and attaching it to a reduced relative clause), but it can also be evoked by semantic violations (see Paczynski and Kuperberg [33] for an example).
@@
Currently, it is unknown whether the confusion induced by atoms of confusion, as perceived by programmers, results (1) from the semantic integration of unexpected and challenging content (in which case we would expect to see an N400), (2) from syntactic ambiguity (see below), or (3) from something else entirely (producing an unexpected ERP signal). As for syntactic ambiguity, the confusion may arise from (2a) syntactic reanalysis (in which case we would expect to see a late posterior positivity) or from (2b) the integration of unexpected but plausible content into the situation model of the preceding context (in which case we would expect to see a late frontal positivity). Thus, we pose the following research question:
@@
The premise of our work is that building on natural-language studies would help us understand the sources and processing of such confusion, illuminating the intricacies of program comprehension. This, in turn, would be invaluable in better supporting programmers through dedicated programming languages and curricula. For example, if atoms of confusion evoke a perceived syntactic ambiguity, this could be avoided by better design choices made by programming language designers (e.g., by requiring explicit brackets around the bodies of for, while, and if statements); if they rather cause problems with semantic integration, this can be captured in dedicated training procedures within programming curricula and coding guidelines (e.g., defining clear and consistent rules that instruct programmers on the use of patterns). If we observe an entirely unexpected ERP signal, this could indicate that the confusion arising from the presence of an atom of confusion has no equivalent in natural-language comprehension, or that there is no unified cause for atoms of confusion. Thus, independent of the concrete results, it is highly relevant for the field of program comprehension to understand the electrophysiological signals and hence the underlying mechanisms of processing atoms of confusion. Looking at a broader picture, if we observe more and more analogies between processing program code and natural language text, this suggests that our premise of a methodological transfer between the two disciplines is worthwhile.
@@
Since it is unclear which ERP component relates to processing confusing code, we employ a data-driven approach to detect differences between confusing and corresponding clean code without targeting a hypothesized component, as research on program comprehension and its underlying neurocognitive mechanisms is still in its infancy. Contrary to research on natural language comprehension, which builds on several decades of behavioral, neurophysiological, and brain imaging studies that have unraveled a variety of processing aspects and representational structures [37], studying program comprehension faces unique challenges, such as a non-linear reading flow (e.g., program statements are processed not in lexical order but along the control and data flow of possible program executions) [38, 39, 40, 41]. So far, only one study on program comprehension has investigated ERP responses to syntactic and semantic errors in program code: Using a token-wise presentation of code, Kuo and Prat [42] identified P600 and N400 components for syntactically and semantically invalid program code, respectively. However, this setup of a token-wise presentation, while usually applied in reading studies with natural language materials [43], is too artificial and highly problematic for the analysis of program code: First, it enforces an unnatural linear reading flow. Second, with a token-wise presentation mode, each token is presented with a uniform duration, which allows the programmer only a cursory glance at the code but does not account for calculations that can strongly vary in complexity between tokens within a code snippet. Both aspects strongly compromise the construct and ecological validity of neurocognitive studies on program comprehension.
@@
In an effort to increase the ecological validity of our study with regard to code presentation and code reading, we seek to present program code as a whole and allowed participants to read the code at their own pace. A major challenge in presenting a snippet of code all at once, however, is that the exact point in time at which the critical segment containing the atom of confusion is processed varies both between items and between participants and therefore cannot be assessed with conventional ERP averaging methods. That is, simply synchronizing electroencephalographical (EEG) activity with the onset of the presentation of the code snippet leads to a temporal misalignment of the ERP activity related to the processing of the atom within the trial period and, thus, diffuses the signal. This severely reduces statistical power and interpretability. To address this key challenge, we employ a more sophisticated and powerful approach based on fixation-related potentials (FRPs). In the FRP approach, event synchronization is not realized by external stimuli [44]. Rather, eye movements and EEG are recorded simultaneously while the program code is presented as a whole, as illustrated in Figure 2a. The eye-tracking data are used to assess the participants’ first eye fixations in the area of the atom of confusion (which is well-defined for every atom of confusion). This first fixation is then used for aligning the EEG signals. This way, we are able to determine the exact point in time at which information from the atoms is processed and can use it for synchronizing the electrophysiological data.
@@
To uncover which neurocognitive mechanisms are engaged in response to processing confusing program code, we conducted an FRP experiment involving participants. We carefully adjusted the existing stimuli from previous work on atoms of confusion in code written in the Java programming language [13] to optimize them for an FRP setup. In particular, each snippet was presented in a condition with and without an atom of confusion (across participants). Each pair of the snippet condition had a comparable number and length of code lines. Furthermore, to select suitable fixations for FRP analysis, we ensured that each snippet contained a single, clearly defined area of interest that included either the atom of confusion (confusing snippets) or an equivalent without an atom (clean snippets) (see Section 4 for details). The experiment consisted of three blocks, as illustrated in Figure 2b, in which we presented snippets (half confusing and half clean). We asked the participants to comprehend the snippets as quickly and accurately as possible in order to report the output. During the experiment, we collected comprehension time, answer correctness, and subjective difficulty ratings to verify that our confusing snippets were generally more difficult to process on the behavioral level than clean snippets. As illustrated in Figure 2a, we collected EEG and eye-tracking data. We used this data to calculate FRPs by mapping fixations to the areas of interest within the code snippets. We took special care to identify and correct eye movements and muscular artifacts by means of independent component analysis (see Section 4, for further details). We synchronized the fixations with the EEG data stream to calculate FRPs. On the obtained FRPs, we tested for significant amplitude differences between confusing and clean code using pairwise cluster permutation tests.
@@ 2 Results
We report on the first-ever FRP analysis of program comprehension by measuring and pinpointing the processing of confusing elements and their neural underpinnings during program comprehension in an ecologically valid setting. Our results show that confusing code elicits a larger frontal positivity between and after first fixating the relevant area of the atom of confusion compared to fixating clean code (without an atom). We relate the identified component to the late frontal positivity occurring in natural language processing for unexpected but plausible input, which elicits an update of the situation model.
@@ 2.1 Behavioral results
The behavioral data comprised comprehension time, answer correctness, and subjective difficulty ratings, where the last two were mapped to a binary categorical scale (the submitted answer was correct or the submitted rating was easy). The descriptive trial statistics are illustrated in the left part of Table 1.
In line with prior work, our behavioral results based on general linear mixed-effects regression (GLMER) models show that the confusing snippets took longer to comprehend, had fewer correct answers, and were rated less frequently as easy to understand, as displayed in the right part of Table 1. The full analysis results, including all fixed and random effects, are provided in the supplementary information, Appendix C.
@@ 2.2 Fixation-related potentials analysis
Our primary goal was to determine the neurocognitive processes triggered by the presence of atoms of confusion in program code and to relate those processes to ERP components typically identified during natural language comprehension. In fact, there is a large delay and large variance in the point in time when participants first fixate on the critical region after the onset of the snippet: the first fixation on the critical region happens, on average, after stimulus onset in the clean condition, and in the confusing condition. The stimulus-onset-aligned ERPs, hence, cannot capture the first point in time at which information from the critical region is presumably processed. Fixation data show that on the first pass, participants fixated the critical region for in the clean condition and in the confusing condition. Therefore, we compared FRPs time-locked to the first fixation in the area of the atom of confusion and its counterpart without atom, spanning an epoch from preceding the start of the fixation until thereafter. For some trials, this epoch could not be extracted due to either low quality in the fixation data, no fixation in the area of interest, artifacts in the EEG signal, or a too short time period until the end of the snippet viewing phase. We excluded these trials from the subsequent analysis ( of ). Then, we calculated the participants’ FRP averages for both conditions across all trials. The grand average FRP data across all participants for electrodes at frontal, central, and parietal recording sites are presented in Figure 3a.
As illustrated in Figure 3, confusing code elicits a sustained frontal positivity relative to clean code. This positivity was most pronounced between and at frontal recording sites (i.e., F3, Fz, F4). At posterior recording sites, the positive amplitude difference between confusing and clean code was attenuated in this time interval. Figure 3b shows the difference in amplitude for all electrodes in the form of topographical distributions in consecutive time intervals between and . Positive differences between both conditions were also present beyond after fixation onset at some left frontal recording sites (e.g., F7 and F3), though they were less consistent. We statistically analyzed these effects using a cluster-based paired permutation t-test, which identifies significant clusters of adjacent data points (defined temporally and spatially).
@@
These results show that the late frontal positivity is significantly greater for confusing code than for clean code, indicating that both conditions are processed differently starting at around after the onset of the first fixation.